OpenStudy (anonymous):

Write an equation in slope-intercept form for the line that satisfies the following condition. passes through (13, 9) and (30, 26

OpenStudy (anonymous):

you need a point and a slope. So to get the slope do (y2-y1)/(x2-x1), then choose either point to use. Then plug into point-slope form and solve for y. y-y1=m(x-x1) so: (26-9)/(30-13)=1 and y-9=1(x-13) solving for y, the answer is y=x-4 Check by plugging your given points back in.